OpenStudy (anonymous):

V=LWH 17=9*6*2 Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the length, then the expression for the height will be (x + 4). What is this asking?

OpenStudy (anonymous):

its looking for the answer of 9 * 6 * 2

OpenStudy (anonymous):

then you have to factor it all out

OpenStudy (anonymous):

its 17

OpenStudy (anonymous):

x * (x - 2) * (x - 8) = 108 i think you have it from here

OpenStudy (anonymous):

no you have to multiply them out not add

OpenStudy (anonymous):

oh, yea wow, im dumb. idk why i added them

OpenStudy (anonymous):

what would be the expression for the height?

OpenStudy (phi):

it sounds like there is more to this question that you left out.

OpenStudy (anonymous):

it does not make sense

OpenStudy (anonymous):

Measure and record the length, width and height of the rectangular box you have chosen in inches. Round to the nearest whole number. Apply the formula of a rectangular box (V = lwh) to find the volume of the object. Now suppose you knew the volume of this object and the relation of the length to the width and height, but did not know the length. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the length. Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the length, then the expression for the height will be (x + 4). This is the whole question. But I already have the L,W & H

OpenStudy (anonymous):

It looks like what they want you to do is take the formula \(V=lwh \) and replace it with \(V=(x+c)(6)(2)\).

OpenStudy (anonymous):

what is c?

OpenStudy (anonymous):

oops! Nevermind that, I think I misread, give me a moment.

OpenStudy (anonymous):

Okay so \(x = l = 9\). This means \(w = x - 3\) and \(h = x-7\) since \(9 - 3 = 6\) and \(9 - 7 = 2\). Make sense so far? I think they then want you to multiply them: \[V = (x)(x-3)(x-7)\]

OpenStudy (anonymous):

how did you get 3?

OpenStudy (phi):

You have length= 9 width=6 height=2 they want you to call the length x they want width= 6 = x+ something in other words 6= 9 + s s= 6-9= -3 so width is (x-3) (where x is 9)

OpenStudy (anonymous):

got it. but then its asking me to simplify the equation and write it in standard form. what equation would that be?

OpenStudy (phi):

before going on, you should fix your original calculation V= 9*6*2 the volume is length TIMES width TIMES height

OpenStudy (anonymous):

that would be 108

OpenStudy (phi):

yes. So you know the answer is 9, 6 and 2 but they want you to pretend you did not know the length is 9, and call it x and you know the width is x-3 (9-3 is 6) and the height is x-7 (9-7 is 2) and when you multiply them you get the volume: x(x-3)(x-7)= 108 standard form is ax^3 + bx^2+cx + d=0 where a,b,c,d are numbers. you have to multiply out your x terms. Can you do that?

OpenStudy (phi):

I would use FOIL to multiply (x-3)*(x-7) first

OpenStudy (anonymous):

(x^2-7x-3x+21)

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

-7x and -3x are like terms.

OpenStudy (phi):

yes, but simplify it -7x-3x = -10 x now multiply it by x

OpenStudy (phi):

x(x-3)(x-7)= 108 x(x^2-10x+21)=108 now mult by x then add -108 to both sides of the = sign

OpenStudy (anonymous):

what do you mean multiply by x, you mean 9?

OpenStudy (anonymous):

is -7x-3x = -10 x, the simplified form?

OpenStudy (phi):

no, not by 9 (we are pretending we don't know what number x is) we are doing x(x-3)(x-7)= 108 here you multiplied (x-3)(x-7) x(x^2-10x+21)=108 now distribute the x (mult every term inside the parens)

OpenStudy (phi):

and yes (x^2-7x-3x+21) is x^2 -10x +21 (after simplifying -7x-3x to be -10x)

OpenStudy (phi):

Here is an example: x(x^2 +2)= x*x^2 +x*2 = x^3 + 2x

OpenStudy (anonymous):

hold on sorry, im gonna read this carefully

OpenStudy (anonymous):

would it be: x^3-10x^2+21x=108 ?

OpenStudy (phi):

yes, one last step to put it in standard form standard form is ax^3 + bx^2+cx + d=0 add -180 to both sides

OpenStudy (anonymous):

what do you mean?

OpenStudy (phi):

you want standard form: stuff on the left side = ZERO you have: stuff on the left= 108 you want to "move" the 108 to the other side of the = sign. write down -108 on both sides of the = sign like this x^3-10x^2+21x -108 =108-108 and simplify

OpenStudy (anonymous):

would i add 108 to the left side?, im sorry i dont get this

OpenStudy (anonymous):

i mean, add it to the right side

OpenStudy (phi):

add -108 (if I typed 180 up above, that is supposed to be 108) to both sides yes. you start with x^3-10x^2+21x=108 you add -108 to BOTH sides x^3-10x^2+21x -108 =108-108

OpenStudy (phi):

This exercise is supposed to help you understand about equations http://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/e/one_step_equation_intuition

OpenStudy (anonymous):

what both sides? i added it to the -108 before the =sign and i got -216 but im assuming thats wrong

OpenStudy (phi):

you start with x^3-10x^2+21x=108 you add -108 to BOTH sides x^3-10x^2+21x -108 =108-108

OpenStudy (anonymous):

ohhhhh, your not telling me to add the -108 just place them their?

OpenStudy (phi):

yes , to both sides (to keep things in balance or equal)

OpenStudy (anonymous):

so that would be standard form?

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

im so confused on what to solve, im not sure how to solve it

OpenStudy (phi):

you are practically done you have x^3-10x^2+21x -108 =108-108 on the right side 108-108 is 0, and this simplifies to x^3-10x^2+21x -108 = 0 that is the polynomial they wanted you to create. If you solve it (find the x that makes the left side =0), you would find x=9 See http://www.wolframalpha.com/input/?i=x%5E3-10x%5E2%2B21x-108%3D0

OpenStudy (anonymous):

which one is the standard form?

OpenStudy (phi):

standard form is ax^3 + bx^2+cx + d=0 can you see that x^3-10x^2+21x -108 = 0 is in this form?

OpenStudy (anonymous):

yes, so i dont have to replace the a,b and c?

OpenStudy (phi):

no, when I write ax^3 + bx^2+cx + d=0, that is just to show you what the equation should look like (we assume a,b,c,d are numbers) See http://www.mathsisfun.com/algebra/standard-form.html (scroll down to polynomial) for more info. But your answer is x^3-10x^2+21x -108 = 0

OpenStudy (anonymous):

awesome, thanks!!!

OpenStudy (phi):

notice it is easier to create this problem (though it did take time), than to solve it. Remember, you want to be the person that creates problems, not the one that has to solve them